OpenStudy (anonymous):

Define Variables, write a system, solve, and answer question. Ben is in charge of ordering meals for a flight on a jet plane that carries first class and coach passengers. Ben spends $5.50 on coach meals for passengers and $9.00 on meals for first class passengers. There are a total of 84 passengers on the plane. If he spends $493.50 all together, how many of each type of passenger is on the plane?

ganeshie8 (ganeshie8):

say, x = coach passengers y = first class passengers

ganeshie8 (ganeshie8):

There are a total of 84 passengers on the plane. that gives, x + y = 84 -----------------(1)

ganeshie8 (ganeshie8):

If he spends $493.50 all together that gives, 5.5x + 9y = 494.5 --------------------(2)

ganeshie8 (ganeshie8):

you have 2 equations, 2 unknowns. you can solve easily

OpenStudy (anonymous):

@ganeshie8 I still don't know what to do from there

ganeshie8 (ganeshie8):

do u knw substitution method ?

OpenStudy (anonymous):

@ganeshie8 Yes but how am i supposed to substitute? I already tried x=84-y but I only get a negative number and that doesn't make sense when I answer the question

ganeshie8 (ganeshie8):

you pluggin in x = 84-y, in equation and you getting negative number for y is it ? lets see

OpenStudy (anonymous):

@ganeshie I'm getting a positive number for y and a negative number for x

ganeshie8 (ganeshie8):

5.5x + 9y = 494.5 5.5(84-y) + 9y = 494.5 462 - 5.5y + 9y = 494.5 3.5y = 32.5 y = 9.286

ganeshie8 (ganeshie8):

ok, thats what u got for y ?

OpenStudy (anonymous):

@ganeshie8 It's 493.5 not 494.5

ganeshie8 (ganeshie8):

ok, what did u get for y ?

ganeshie8 (ganeshie8):

thanks for correcting :)

OpenStudy (anonymous):

@ganeshie8 I got 9

ganeshie8 (ganeshie8):

5.5x + 9y = 493.5 5.5(84-y) + 9y = 493.5 462 - 5.5y + 9y = 493.5 3.5y = 31.5 y = 9

ganeshie8 (ganeshie8):

good :) i got the same

OpenStudy (anonymous):

@ganeshie8 and x is 75?

ganeshie8 (ganeshie8):

Yes ! very good work :D

OpenStudy (anonymous):

@ganeshie8 Thank You! :)

ganeshie8 (ganeshie8):

np :)